# From tug-of-war to Brownian Boost: explicit ODE solutions for player-funded stochastic-differential games

### v3

@@ 4.5.2. Overview of the proof at large
Given the reduction of the proof of Theorem 1.8(2) that we summarised verbally after Propositions 4.16 and 4.17, which is recorded more formally in [26, Section $4.2$], the substantial elements for this overview are the proofs of this pair of results. We discuss them in turn.
Deriving Proposition 4.16. There may be infinitely many Mina deviation points for $S_{-}^{\mathrm{dev}}$ whose spatial coordinate lies in $\llbracket-j,k\rrbracket$. We begin by reducing to a finite number by eliminating late deviating moves. For $h\in\mathbb{N}$, let the strategy $S_{-}^{\mathrm{dev}}[h]$ be formed from $S_{-}^{\mathrm{dev}}$ by removing every deviating move after time $h$: thus, Mina will stake $b_{u}$ at $(u,t)\in\mathbb{Z}\times\mathbb{N}$ when $t\geq h$. Since $\mathbb{P}_{S_{-}^{\mathrm{dev}},S_{+}}^{i}(E)=1$, the strategy pair $(S_{-}^{\mathrm{dev}},S_{+})$ when played from $i$ on gameboard $\llbracket-j,k\rrbracket$ results in termination at a random finite time; so if Mina plays $S_{-}^{\mathrm{dev}}[h]$ in place of $S_{-}^{\mathrm{dev}}$ for high $h$, there will be merely an arbitrarily small shift in the mean outcomes.
@@
To state the formal change needed: the two displayed equations in the proof of [26, Lemma 4.16(2)] will now read
$\mathbb{E}^{u,\ell}_{S,S_{+}}\big[P^{j,k}_{-}\big]$ $=$ $\bigg(\frac{\kappa\,S(u,\ell)^{\rho}}{a_{i}^{\rho}+S(u,\ell)^{\rho}}+\frac{1-\kappa}{2}\bigg)\,\mathbb{E}^{u-1,\ell+1}_{S,S_{+}}\big[P^{j,k}_{-}\big]$
$\qquad\qquad\qquad+\,\bigg(\frac{\kappa\,a_{i}^{\rho}}{a_{i}^{\rho}+S(u,\ell)^{\rho}}+\frac{1-\kappa}{2}\bigg)\,\mathbb{E}^{u+1,\ell+1}_{S,S_{+}}\big[P^{j,k}_{-}\big]\,-\,S(u,\ell)$
(with the superscript in the notation $\mathbb{E}^{u\pm 1,\ell+1}_{S,S_{+}}$ set out in [26, Section $3.3$] referring to delayed-start games); and
$\mathbb{E}^{u,\ell}_{S,S_{+}}\big[P^{j,k}_{-}\big]\,\leq\,\bigg(\frac{\kappa\,S(u,\ell)^{\rho}}{a_{i}^{\rho}+S(u,\ell)^{\rho}}\,+\,\frac{1-\kappa}{2}\bigg)n_{u-1}+\bigg(\frac{\kappa\,a_{i}^{\rho}}{a_{i}^{\rho}+S(u,\ell)^{\rho}}+\frac{1-\kappa}{2}\bigg)n_{u+1}\,-\,S(u,\ell)\,.$
@@
A second variation addresses a point that has been elided in the above summary. There is a difference in strategy definition between [26] and the present article. While [26] specifies strategies simply as functions of space-time, we permit them to depend on the counter history to the present moment. This has led us to a definition of Mina deviation point whereby there must exist at least one history leading to the point in question which would cause her to place a deviant stake in playing from there. In order that the proof of Proposition 4.16 leads to a strict inequality in its conclusion, it is enough to argue that, for at least one Mina deviation point $(u,\ell)$ with $u\in\llbracket-j,k\rrbracket$, every element in the path space $\Lambda$ that begins at $(i,0)$ and ends at $(u,\ell)$ lies in the rectangle $\llbracket-j,k\rrbracket\times\llbracket 0,\ell\rrbracket$. Indeed, for such a point $(u,\ell)$, there exists a history $(0,0)\to(u,\ell)$—choose one and call it hh!—which induces Mina to play a deviant move at the $(\ell+1)$st turn. The counter may follow this path without the game played on $\llbracket-j,k\rrbracket$ ending. The trajectory follows this history with positive probability (if $\kappa\in(0,1)$, via a sequence of flip moves; if $\kappa=1$, by an argument in [26]). Consequently, the introduction of this Mina deviation point in the iterative procedure discussed above leads to a positive loss in her mean payoff, as in the proof we are adapting. The loss is determined by $(u,\ell)$ and $h$. The introduction of other deviation points has a non-positive effect on her payoff, so the cumulative effect is bounded above by the said loss. In Proposition 4.16, a given deviation point $(u,\ell)$ is considered, and the hypothesis $u\in\llbracket-j+\ell,k$-$\ell\rrbracket$ is imposed on $j$ and $k$. It is this hypothesis that ensures that $(u,\ell)$ meets the condition on path-space inclusion. The values of $j$ and $k$ may be chosen to exceed some large constant specified by the given $(u,\ell)$, so the resulting loss is independent of such $(j,k)$; this leads to the uniformity asserted in Proposition 4.16.
@@
In summary, an inductive argument based on noting that deviating play is punished in the one-step game leads to the inference that the above discussed finite-deviating strategies $S_{-}^{\mathrm{dev}}[h]$ are uniformly punished on finite trails. By choosing the finite trails to be broad enough, the condition $\mathbb{P}_{S_{-}^{\mathrm{dev}},S_{+}}^{i}(E)=1$ implies that the error arising from the use of $S_{-}^{\mathrm{dev}}[h]$ in place of $S_{-}^{\mathrm{dev}}$ is for high $h$ smaller than the incurred penalty. In this way, Proposition 4.16.is derived.
@@
Obtaining Proposition 4.17 from Proposition 4.16. Proposition 4.18 permits us to reduce to the case where $\mathbb{P}^{i}_{S_{-}^{\mathrm{dev}},S_{+}}(E)=1$. The certainty of escape means that the counter will leave a broad enough board on the side on which it escapes globally. This permits us to truncate to a broad finite board incurring an arbitrarily small discrepancy in mean terminal payment. Removing non-negative running costs incurred beyond departure from the finite board then yields $\mathbb{E}_{S_{-}^{\mathrm{dev}},S_{+}}^{i}[P_{-}]\leq\mathbb{E}_{S_{-}^{\mathrm{dev}},S_{+}}^{i}[P_{-}^{j,k}]$ up to the same small error. Proposition 4.16 may then be invoked, with the uniform penalty there identified overcoming the small opposing error, yielding the sought bound $\mathbb{E}_{S_{-}^{\mathrm{dev}},S_{+}}^{i}[P_{-}]<\mathbb{E}_{S_{-},S_{+}}^{i}[P_{-}]$.
@@
Proof of Proposition 4.18. The trap event $\mathsf{Trap}$ is a costly one for Mina because her terminal receipt in this event will be $n_{*}$, which is by assumption strictly lower than her losing receipt $n_{\infty}$; and, moreover, she may have running costs to pay. She would be happier with an altered strategy in which she instead consistently stakes zero in the trapping event, leading to an improved terminal receipt of $n_{\infty}$ alongside zero running cost. The problem with this idea is that the proposed alteration is not a well-defined strategy, because the proposed change is contingent on the occurrence of $E^{c}=\mathsf{Trap}$, an event undetermined by any finite-step evaluation of gameplay. We will resolve this difficulty by introducing an event $\mathsf{ProxyTrap}$ determined by an initial move-sequence that nearly coincides with $\mathsf{Trap}$, and defining Mina’s altered strategy $S_{-}^{\mathrm{dev}}[\mathrm{alt}]$ by asking her to stake zero after the moment at which $\mathsf{ProxyTrap}$ has been determined to occur. The definition will yield an admissible strategy because the specification of the strategy space $\mathcal{S}$ in Section 1.3 permits a player to consult counter history in deciding how to stake.
@@
Now assume that $\mathbb{P}_{S_{-}^{\mathrm{dev}},S_{+}}^{i}(\mathsf{Trap})\in(0,1)$. In constructing and analysing the altered strategy $S_{-}^{\mathrm{dev}}[\mathrm{alt}]$, we will couple the gameplays under $(S_{-}^{\mathrm{dev}},S_{+})$ and $(S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+})$. We will write $\mathbb{P}^{i}$ for the law governing these two gameplays, and will distinguish between them by indicating the strategy pair associated to a given random variable. For example, $T_{-}(S_{-}^{\mathrm{dev}},S_{+})$ under $\mathbb{P}^{i}$ denotes Mina’s terminal payoff for gameplay governed by the strategy pair $(S_{-}^{\mathrm{dev}},S_{+})$ under the coupling; in law, this random variable is equal to $T_{-}$ under $\mathbb{P}^{i}_{S_{-}^{\mathrm{dev}},S_{+}}$.
Under $\mathbb{P}^{i}$, we will define $\mathsf{ProxyTrap}$ in terms of a parameter $\epsilon>0$ measuring the approximation of $\mathsf{Trap}$. We will set $\mathsf{ProxyTrap}=\big\{\tau_{\epsilon}<\infty\big\}$ for an $\mathbb{N}$-valued stopping time $\tau_{\epsilon}$ in such a way that
$\textrm{$\mathsf{Trap}\subseteq\mathsf{ProxyTrap}$ holds up to a $\mathbb{P}^{i}$-null set, and}\,\,\,\,\mathbb{P}^{i}\big(\mathsf{ProxyTrap}\setminus\mathsf{Trap}\big)\leq\epsilon\,.$
To construct $\tau_{\epsilon}$, let $_{j}$ for $j\in\mathbb{Z}$ denote the total number of visits made by $X:\mathbb{N}\to\mathbb{Z}$ to the site $j$, for the copy of counter evolution under $(S_{-}^{\mathrm{dev}},S_{+})$ offered by $\mathbb{P}^{i}$. It follows readily from the meaning of absence of escape that for any $i\in\mathbb{N}_{+}$, we may find a non-random finite subset $J_{i}\subset\mathbb{Z}$ such that
$\mathbb{P}^{i}\Big(\max_{j\in J_{i}}\#_{j}=\infty\,\Big|\,\mathsf{Trap}\Big)\,\geq\,1-\epsilon/2^{i}\,$, (38)
@@
$\mathbb{P}^{i}\Big(\max_{j\in J_{i}}\#_{j}\geq N_{i}\,\Big|\,\mathsf{Trap}^{c}\Big)\,<\,\epsilon/2^{i}\,.$ (39)
@@
Now we set $\tau_{\epsilon}=\min_{i\in\mathbb{N}}\phi_{i}$. To define $S_{-}^{\mathrm{dev}}[\mathrm{alt}]$, recall the path spaces $\Lambda_{k}$ used to specify $\mathcal{S}$ in Section 1.3. We set $S_{-}^{\mathrm{dev}}[\mathrm{alt}](\psi)=S_{-}^{\mathrm{dev}}(\psi)$ whenever $\psi\in\Lambda_{k}$ for some $k\in\mathbb{N}$ such that $\tau_{\epsilon}>k$ for any counter evolution $X$ with $X\big|_{\llbracket 0,k\rrbracket}=\psi$. The value of $S_{-}^{\mathrm{dev}}[\mathrm{alt}](\psi)$ is set to zero for any $\psi$ in the remainder of $\Lambda$. It is a straightforward check that $\tau_{\epsilon}$ is a stopping time and $S_{-}^{\mathrm{dev}}[\mathrm{alt}]$ an element of $\mathcal{S}$. The coupling $\mathbb{P}^{i}$ is constructed so that counter evolutions under $(S_{-}^{\mathrm{dev}},S_{+})$ and $(S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+})$ almost surely coincide until  $\tau_{\epsilon}$.
Next we verify the desired properties that $\mathsf{Trap}\subset\mathsf{ProxyTrap}$ and $\mathbb{P}^{i}(\mathsf{ProxyTrap}\setminus\mathsf{Trap})<\epsilon$. To do so, note that (38) implies that, conditionally on $\mathsf{Trap}$, $\max_{j\in J_{i}}\#_{j}$ equals infinity for all but finitely many $i$ almost surely, so that $\phi_{i}$ is finite with the exception of at most finitely many $i$; this implies that $\tau_{\epsilon}<\infty$, so that $\mathsf{ProxyTrap}$ occurs. Note further that
$\mathbb{P}^{i}(\mathsf{ProxyTrap}\setminus\mathsf{Trap})=\mathbb{P}\Big(\mathsf{Trap}^{c}\cap\big\{\exists\,i\in\mathbb{N}_{+},j\in J_{i}:\#_{j}\geq N_{i}\big\}\Big)<\sum_{i=1}^{\infty}\ \epsilon/2^{i}=\epsilon\,$,
the bound due to (39).
We now use the constructed $\tau_{\epsilon}$ to prove that the desired (37) holds. Note that
$\mathbb{E}^{i}\Big[P_{-}(S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+}){\bf 1}_{\mathsf{ProxyTrap}^{c}}\Big]=\mathbb{E}^{i}\Big[P_{-}(S_{-}^{\mathrm{dev}},S_{+}){\bf 1}_{\mathsf{ProxyTrap}^{c}}\Big]\,.$ (40)
@@
$\mathbb{E}^{i}\Big[P_{-}(S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+}){\bf 1}_{\mathsf{Trap}}\Big]$ $=$ $\mathbb{E}^{i}\,\bigg[\Big(T_{-}(S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+})-\sum_{i=0}^{\infty}S_{-}^{\mathrm{dev}}[\mathrm{alt}](i)\Big){\bf 1}_{\mathsf{Trap}}\bigg]$
$\geq$ $\big(n_{\infty}-n_{*}\big)\mathbb{P}^{i}(\mathsf{Trap})\,+\,\mathbb{E}^{i}\,\bigg[\Big(T_{-}(S_{-}^{\mathrm{dev}},S_{+})-\sum_{i=0}^{\infty}S_{-}^{\mathrm{dev}}(i)\Big){\bf 1}_{\mathsf{Trap}}\bigg]$
$=$ $\big(n_{\infty}-n_{*}\big)\mathbb{P}_{(S_{-}^{\mathrm{dev}},S_{+})}(\mathsf{Trap})+\mathbb{E}_{S_{-}^{\mathrm{dev}},S_{+}}[P_{-}{\bf 1}_{\mathsf{Trap}}]\,.$
@@
$\mathbb{E}^{i}\Big[P_{-}(S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+}){\bf 1}_{\mathsf{ProxyTrap}\setminus\mathsf{Trap}}\Big]-\mathbb{E}^{i}\Big[P_{-}(S_{-}^{\mathrm{dev}},S_{+}){\bf 1}_{\mathsf{ProxyTrap}\setminus\mathsf{Trap}}\Big]$
$\geq$ $\mathbb{E}^{i}\Big[T_{-}(S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+}){\bf 1}_{\mathsf{ProxyTrap}\setminus\mathsf{Trap}}\Big]-\mathbb{E}^{i}\Big[T_{-}(S_{-}^{\mathrm{dev}},S_{+}){\bf 1}_{\mathsf{ProxyTrap}\setminus\mathsf{Trap}}\Big]$
$\geq$ $\Big(n_{\infty}-\big(\mu n_{\infty}+(1-\mu)n_{-\infty}\big)\Big)\mathbb{P}^{i}\big(\mathsf{ProxyTrap}\setminus\mathsf{Trap}\big)\geq-\big(n_{\infty,-\infty}\big)\epsilon\,$,

### v4

@@ 4.5.2. Overview of the proof at large
Given the reduction of the proof of Theorem 1.3(2) that we summarised verbally after Propositions 4.16 and 4.17, which is recorded more formally in [24, Section $4.2$], the substantial elements for this overview are the proofs of this pair of results. We discuss them in turn.
Deriving Proposition 4.16. There may be infinitely many Mina deviation points for $S_{-}^{\mathrm{dev}}$ whose spatial coordinate lies in $\llbracket-j,k\rrbracket$. We begin by reducing to a finite number by eliminating late deviating moves. For $R\in\mathbb{N}$, let the strategy $S_{-}^{\mathrm{dev}}[R]$ be formed from $S_{-}^{\mathrm{dev}}$ by removing every deviating move after time $R$: thus, Mina will stake $b_{u}$ at $(u,t)\in\mathbb{Z}\times\mathbb{N}$ when $t\geq R$. Since $\mathbb{P}_{S_{-}^{\mathrm{dev}},S_{+}}^{h}(E)=1$, the strategy pair $(S_{-}^{\mathrm{dev}}[R],S_{+})$ when played from $h$ on gameboard $\llbracket-j,k\rrbracket$ results in termination at a random finite time; so if Mina plays $S_{-}^{\mathrm{dev}}[R]$ in place of $S_{-}^{\mathrm{dev}}$ for high $R$, there will be merely an arbitrarily small shift in the mean outcomes.
@@
To state the formal change needed: we write $\mathbb{P}_{S,S_{+}}^{u,\ell}$ for the law of a delayed-start game (governed by the indicated strategy pair), in accordance with notation set out in [24, Section $3.3$]. Under this law, the counter starts at $u$ on the turn with index $\ell$. We record the stake offered under strategy $S$ at this starting turn as $S(u,\ell)$, with the history that brought the counter to this space-time point being implicit. The two displayed equations in the proof of [24, Lemma 4.16(2)] will now read
$\mathbb{E}^{u,\ell}_{S,S_{+}}\big[P^{j,k}_{-}\big]$ $=$ $\bigg(\frac{\kappa\,S(u,\ell)^{\rho}}{a_{u}^{\rho}+S(u,\ell)^{\rho}}+\frac{1-\kappa}{2}\bigg)\,\mathbb{E}^{u-1,\ell+1}_{S,S_{+}}\big[P^{j,k}_{-}\big]$
$\qquad\qquad\qquad+\,\bigg(\frac{\kappa\,a_{u}^{\rho}}{a_{u}^{\rho}+S(u,\ell)^{\rho}}+\frac{1-\kappa}{2}\bigg)\,\mathbb{E}^{u+1,\ell+1}_{S,S_{+}}\big[P^{j,k}_{-}\big]\,-\,S(u,\ell)\,$,
and
$\mathbb{E}^{u,\ell}_{S,S_{+}}\big[P^{j,k}_{-}\big]\,\leq\,\bigg(\frac{\kappa\,S(u,\ell)^{\rho}}{a_{u}^{\rho}+S(u,\ell)^{\rho}}\,+\,\frac{1-\kappa}{2}\bigg)n_{u-1}+\bigg(\frac{\kappa\,a_{u}^{\rho}}{a_{u}^{\rho}+S(u,\ell)^{\rho}}+\frac{1-\kappa}{2}\bigg)n_{u+1}\,-\,S(u,\ell)\,.$
@@
A second variation addresses a point that has been elided in the above summary. There is a difference in strategy definition between [24] and the present article. While [24] specifies strategies simply as functions of space-time, we permit them to depend on the counter history to the present moment. This has led us to a definition of Mina deviation point whereby there must exist at least one history leading to the point in question which would cause her to place a deviant stake in playing from there. In order that the proof of Proposition 4.16 leads to a strict inequality in its conclusion, it is enough to argue that, for at least one Mina deviation point $(u,\ell)$ with $u\in\llbracket-j,k\rrbracket$, every element in the path space $\Lambda$ that begins at $(h,0)$ and ends at $(u,\ell)$ lies in the rectangle $\llbracket-j,k\rrbracket\times\llbracket 0,\ell\rrbracket$. Indeed, for such a point $(u,\ell)$, there exists a history $(h,0)\to(u,\ell)$—choose one and call it HH!—which induces Mina to play a deviant move at the $(\ell+1)$st turn. The counter may follow this path without the game played on $\llbracket-j,k\rrbracket$ ending. The trajectory follows this history with positive probability (if $\kappa\in(0,1)$, via a sequence of flip moves; if $\kappa=1$, by an argument in [24]). Consequently, the introduction of this Mina deviation point in the iterative procedure discussed above leads to a positive loss in her mean payoff, as in the proof we are adapting. The loss is determined by $(u,\ell)$ and $H$. The introduction of other deviation points has a non-positive effect on her payoff, so the cumulative effect is bounded above by the said loss. In Proposition 4.16, a given deviation point $(u,\ell)$ is considered, and the hypothesis $u\in\llbracket-j+\ell,k$-$\ell\rrbracket$ is imposed on $j$ and $k$. It is this hypothesis that ensures that $(u,\ell)$ meets the condition on path-space inclusion. The values of $j$ and $k$ may be chosen to exceed some large constant specified by the given $(u,\ell)$, so the resulting loss is independent of such $(j,k)$; this leads to the uniformity asserted in Proposition 4.16.
@@
In summary, an inductive argument based on noting that deviating play is punished in the one-step game leads to the inference that the above discussed finite-deviating strategies $S_{-}^{\mathrm{dev}}[R]$ are uniformly punished on finite trails. By choosing the finite trails to be broad enough, the condition $\mathbb{P}_{S_{-}^{\mathrm{dev}},S_{+}}^{h}(E)=1$ implies that the error arising from the use of $S_{-}^{\mathrm{dev}}[R]$ in place of $S_{-}^{\mathrm{dev}}$ is for high $R$ smaller than the incurred penalty. In this way, Proposition 4.16 is derived.
@@
Obtaining Proposition 4.17 from Proposition 4.16. Proposition 4.18 permits us to reduce to the case where $\mathbb{P}^{h}_{S_{-}^{\mathrm{dev}},S_{+}}(E)=1$. The certainty of escape means that the counter will leave a broad enough board on the side on which it escapes globally. This permits us to truncate to a broad finite board incurring an arbitrarily small discrepancy in mean terminal payment. Removing non-negative running costs incurred beyond departure from the finite board then yields $\mathbb{E}_{S_{-}^{\mathrm{dev}},S_{+}}^{h}[P_{-}]\leq\mathbb{E}_{S_{-}^{\mathrm{dev}},S_{+}}^{h}[P_{-}^{j,k}]$ up to the same small error. Proposition 4.16 may then be invoked, with the uniform penalty there identified overcoming the small opposing error, yielding the sought bound $\mathbb{E}_{S_{-}^{\mathrm{dev}},S_{+}}^{h}[P_{-}]<\mathbb{E}_{S_{-},S_{+}}^{h}[P_{-}]$.
@@
Proof of Proposition 4.18. The trap event $\mathsf{Trap}$ is a costly one for Mina because her terminal receipt in this event will be $n_{*}$, which is by assumption strictly lower than her losing receipt $n_{\infty}$; and, moreover, she may have running costs to pay. She would be happier with an altered strategy in which she instead consistently stakes zero in the trapping event, leading to an improved terminal receipt of $n_{\infty}$ alongside zero running cost. The problem with this idea is that the proposed alteration is not a well-defined strategy, because the proposed change is contingent on the occurrence of $E^{c}=\mathsf{Trap}$, an event undetermined by any finite-step evaluation of gameplay. We will resolve this difficulty by introducing an event $\mathsf{ProxyTrap}$ determined by an initial move-sequence that nearly coincides with $\mathsf{Trap}$, and defining Mina’s altered strategy $S_{-}^{\mathrm{dev}}[\mathrm{alt}]$ by asking her to stake zero after the moment at which $\mathsf{ProxyTrap}$ has been determined to occur. The definition will yield an admissible strategy because the specification of the strategy space $\mathcal{S}$ in Section 1.3.1 permits a player to consult counter history in deciding how to stake.
@@
Now assume that $\mathbb{P}_{S_{-}^{\mathrm{dev}},S_{+}}^{h}(\mathsf{Trap})\in(0,1)$. In constructing and analysing the altered strategy $S_{-}^{\mathrm{dev}}[\mathrm{alt}]$, we will couple the gameplays under $(S_{-}^{\mathrm{dev}},S_{+})$ and $(S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+})$. We will write $\mathbb{P}^{h}$ for the law governing these two gameplays, and will distinguish between them by indicating the strategy pair associated to a given random variable. For example, $T_{-}(S_{-}^{\mathrm{dev}},S_{+})$ under $\mathbb{P}^{h}$ denotes Mina’s terminal payoff for gameplay governed by the strategy pair $(S_{-}^{\mathrm{dev}},S_{+})$ under the coupling; in law, this random variable is equal to $T_{-}$ under $\mathbb{P}^{h}_{S_{-}^{\mathrm{dev}},S_{+}}$.
Under $\mathbb{P}^{h}$, we will define $\mathsf{ProxyTrap}$ in terms of a parameter $\epsilon>0$ measuring the approximation of $\mathsf{Trap}$. We will set $\mathsf{ProxyTrap}=\big\{\tau_{\epsilon}<\infty\big\}$ for an $\mathbb{N}$-valued stopping time $\tau_{\epsilon}$ in such a way that
$\textrm{$\mathsf{Trap}\subseteq\mathsf{ProxyTrap}$ holds up to a $\mathbb{P}^{h}$-null set, and}\,\,\,\,\mathbb{P}^{h}\big(\mathsf{ProxyTrap}\setminus\mathsf{Trap}\big)\leq\epsilon\,.$
To construct $\tau_{\epsilon}$, let $_{j}$ for $j\in\mathbb{Z}$ denote the total number of visits made by $X:\mathbb{N}\to\mathbb{Z}$ to the site $j$, for the copy of counter evolution under $(S_{-}^{\mathrm{dev}},S_{+})$ offered by $\mathbb{P}^{h}$. It follows readily from the meaning of absence of escape that for any $i\in\mathbb{N}_{+}$, we may find a non-random finite subset $J_{i}\subset\mathbb{Z}$ such that
$\mathbb{P}^{h}\Big(\max_{j\in J_{i}}\#_{j}=\infty\,\Big|\,\mathsf{Trap}\Big)\,\geq\,1-\epsilon/2^{i}\,$, (4.12)
@@
$\mathbb{P}^{h}\Big(\max_{j\in J_{i}}\#_{j}\geq N_{i}\,\Big|\,\mathsf{Trap}^{c}\Big)\,<\,\epsilon/2^{i}\,.$ (4.13)
@@
Now we set $\tau_{\epsilon}=\min_{i\in\mathbb{N}}\phi_{i}$. To define $S_{-}^{\mathrm{dev}}[\mathrm{alt}]$, recall the path spaces $\Lambda_{k}$ used to specify $\mathcal{S}$ in Section 1.3.1. We set $S_{-}^{\mathrm{dev}}[\mathrm{alt}](\psi)=S_{-}^{\mathrm{dev}}(\psi)$ whenever $\psi\in\Lambda_{k}$ for some $k\in\mathbb{N}$ such that $\tau_{\epsilon}>k$ for any counter evolution $X$ with $X\big|_{\llbracket 0,k\rrbracket}=\psi$. The value of $S_{-}^{\mathrm{dev}}[\mathrm{alt}](\psi)$ is set to zero for any $\psi$ in the remainder of $\Lambda$. It is a straightforward check that $\tau_{\epsilon}$ is a stopping time and $S_{-}^{\mathrm{dev}}[\mathrm{alt}]$ an element of $\mathcal{S}$ (which coincides with $S_{-}^{\mathrm{dev}}$ before $\tau_{\epsilon}$ and is zero at that time and thereafter). The coupling $\mathbb{P}^{h}$ is constructed so that counter evolutions under $(S_{-}^{\mathrm{dev}},S_{+})$ and $(S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+})$ almost surely coincide until  $\tau_{\epsilon}$.
Next we verify the desired properties that $\mathsf{Trap}\subset\mathsf{ProxyTrap}$ and $\mathbb{P}^{h}(\mathsf{ProxyTrap}\setminus\mathsf{Trap})<\epsilon$. To do so, note that (4.12) implies that, conditionally on $\mathsf{Trap}$, $\max_{j\in J_{i}}\#_{j}$ equals infinity for all but finitely many $i$ almost surely, so that $\phi_{i}$ is finite with the exception of at most finitely many $i$; this implies that $\tau_{\epsilon}<\infty$, so that $\mathsf{ProxyTrap}$ occurs. Note further that
$\mathbb{P}^{h}(\mathsf{ProxyTrap}\setminus\mathsf{Trap})=\mathbb{P}^{h}\Big(\mathsf{Trap}^{c}\cap\big\{\exists\,i\in\mathbb{N}_{+},j\in J_{i}:\#_{j}\geq N_{i}\big\}\Big)<\sum_{i=1}^{\infty}\ \epsilon/2^{i}=\epsilon\,$,
the bound due to (4.13).
We now use the constructed $\tau_{\epsilon}$ to prove that the desired (4.11) holds. Note that
$\mathbb{E}^{h}\Big[P_{-}(S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+}){\bf 1}_{\mathsf{ProxyTrap}^{c}}\Big]=\mathbb{E}^{h}\Big[P_{-}(S_{-}^{\mathrm{dev}},S_{+}){\bf 1}_{\mathsf{ProxyTrap}^{c}}\Big]\,.$ (4.14)
@@
$\mathbb{E}^{h}\Big[P_{-}(S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+}){\bf 1}_{\mathsf{Trap}}\Big]$ $=$ $\mathbb{E}^{h}\,\bigg[\Big(T_{-}(S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+})-\sum_{i=0}^{\infty}S_{-}^{\mathrm{dev}}[\mathrm{alt}](i)\Big){\bf 1}_{\mathsf{Trap}}\bigg]$
$\geq$ $\big(n_{\infty}-n_{*}\big)\mathbb{P}^{h}(\mathsf{Trap})\,+\,\mathbb{E}^{h}\,\bigg[\Big(T_{-}(S_{-}^{\mathrm{dev}},S_{+})-\sum_{i=0}^{\infty}S_{-}^{\mathrm{dev}}(i)\Big){\bf 1}_{\mathsf{Trap}}\bigg]$
$=$ $\big(n_{\infty}-n_{*}\big)\mathbb{P}^{h}(\mathsf{Trap})+\mathbb{E}^{h}\big[P_{-}(S_{-}^{\mathrm{dev}},S_{+}){\bf 1}_{\mathsf{Trap}}\big]\,.$
@@
$\mathbb{E}^{h}\Big[P_{-}(S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+}){\bf 1}_{\mathsf{ProxyTrap}\setminus\mathsf{Trap}}\Big]-\mathbb{E}^{h}\Big[P_{-}(S_{-}^{\mathrm{dev}},S_{+}){\bf 1}_{\mathsf{ProxyTrap}\setminus\mathsf{Trap}}\Big]$
$\geq$ $\mathbb{E}^{h}\Big[T_{-}(S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+}){\bf 1}_{\mathsf{ProxyTrap}\setminus\mathsf{Trap}}\Big]-\mathbb{E}^{h}\Big[T_{-}(S_{-}^{\mathrm{dev}},S_{+}){\bf 1}_{\mathsf{ProxyTrap}\setminus\mathsf{Trap}}\Big]$
$\geq$ $\Big(n_{\infty}-\big(\mu n_{\infty}+(1-\mu)n_{-\infty}\big)\Big)\mathbb{P}^{h}\big(\mathsf{ProxyTrap}\setminus\mathsf{Trap}\big)\geq-\big(n_{\infty,-\infty}\big)\epsilon\,$,